According to the ministry of health, the height of Guinean travellers who were quarantined in Tamale for the novel coronavirus were normally distributed about a mean of 160cm and a standard deviation of 8cm. Find the probability that a traveller selected at random has
1
Expert's answer
2020-06-04T20:18:45-0400

Let "X=" the height of the traveller:"X\sim N(\mu, \sigma^2)." Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0,1)"

Given "\mu=160\ cm,\sigma=8\ cm"

Find the probability that a traveller selected at random has 

i) Height between 148cm and 175cm


"P(148<X<175)=P(X<175)-P(X\leq 148)="

"=P(Z<{175-160\over 8})-P(Z\leq{148-160\over 8})="

"=P(Z<1.875)-P(Z\leq-1.5)\approx"

"\approx0.9696-0.0668=0.9028"

ii) Height above 164cm


"P(X>164)=1-P(X\leq 164)=1-P(Z\leq{164-160\over 8})="

"=1-P(Z\leq0.5)\approx1-0.6915=0.3085"

iii) Height below 179cm


"P(X<179)=P(Z<{179-160\over 8})=P(Z<2.375)\approx"

"\approx0.9912"